OpenStudy (anonymous):

The sum of two numbers is 4 less than twice the larger. If the larger is decreased by three times the smaller, the result is -20. Find the numbers

OpenStudy (anonymous):

equation: let x be the smaller number y be the larger number (1) x+y = 2y-4 y-3x=-20

OpenStudy (anonymous):

Let the two numbers be x and y. x+y=2x-4(say,x is the larger number). Now by the question, x-3y=-20. Now solve for x and y, -x+y=-4, and x-3y=-20.

OpenStudy (anonymous):

Let x be the lager number. let y be the smaller number then the equations are x + y = 2x -4 x - 3y = -20 Rearange the equations as follows: -x + y = -4 ( after subtracting 2x from both sides) x -3y = -20 add both equations to eliminate x i.e. -2y = -24 y =12 Plug in the value of y in any one of th above equations to yield -x +12 = -4 =>-x = -16 => x = 16 Therefore larger number = 16 Smaller number =12